A coil of wire is situated in a 0.5 T uniform magnetic field. The area of the coil is 2.0 m2. (a) What is the magnetic flux if the angle between the magnetic field and the normal to the surface of the coil is 60°? (b) After 5 s, the magnetic field is now parallel to the normal to the surface, (at 0°).

1
Expert's answer
2022-04-11T12:07:00-0400

(a)

"\Phi_1=BA\cos\theta\\
=0.5*2.0*\cos60^\circ=0.5\:\rm Wb"

(b)

"\Phi_2=BA\cos\theta\\
=0.5*2.0*\cos0^\circ=1.0\:\rm Wb"